OpenStudy (anonymous):

Which of the following is the solution to the equation 25^(z – 2) = 125 ?

OpenStudy (turingtest):

rewrite 25 and 125 as powers of 5, and apply the rule\[\large (x^a)^b=x^{ab}\]then both sides will have the same base, which you can take the logarithm of

OpenStudy (anonymous):

i have no idea how to do it at all

OpenStudy (turingtest):

\[\large25=5^?\]

OpenStudy (turingtest):

i.e. 25 is 5 raised to what power?

OpenStudy (anonymous):

second

OpenStudy (turingtest):

good what about 125 ?

OpenStudy (anonymous):

5^3?

OpenStudy (turingtest):

yep :) so now we have\[\large(5^2)^{z-2}=5^3\]now we use the log rule I stated above\[\large(x^a)^b=x^{ab}\]to get\[\large5^{2z-4}=5^3\]make sense?

OpenStudy (anonymous):

yes but i dont understand how to get the answer. is it -.5?

OpenStudy (turingtest):

no... to get the final answer take the log base 5 of both sides, then solve the resulting (linear) equation for x if this is a novel concept to you then you must have missed a lot...

OpenStudy (anonymous):

yea i was going thru alot at home and missed a lot of school :// so the 2z-4 =3?